Find two consecutive even integers such that the sum of the larger and twice the smaller is 62
Answer by mananth(16946)

Find two consecutive even integers such that the sum of the larger and twice the smaller is 62

Let the even numbers be x & x+2
(x+2)+2x=62
x+2+2x=62
3x+2=62
3x=62-2
3x=60
/3
x=20
20 & 22 are the numbers